OpenStudy (anonymous):

Explain the difference between using the trigonometric ratios (sin, cos, tan) to solve for a missing angle in a right triangle versus using the reciprocal ratios (sec, csc, cot). You must use complete sentences and any evidence needed (such as an example) to prove your point of view. @hero

OpenStudy (anonymous):

Can anyone please help me!!

OpenStudy (anonymous):

@hero

Nnesha (nnesha):

they are reciprocal not much difference  \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]